OpenStudy (lgg23):

Related Rates Problem: A plane flying with a constant speed of 26 km/min passes over a ground radar station at an altitude of 6 km and climbs at an angle of 35 degrees. At what rate, in km/min is the distance from the plane to the radar station increasing 3 minutes later? Hint: The law of cosines for a triangle is c^2 = a^2 + b^2 - 2ab cos(theta) where theta is the angle between the sides of length a and b.

OpenStudy (anonymous):

do you have a picture?

OpenStudy (anonymous):



OpenStudy (anonymous):

how is that. the 145 from 180 - 35 degrees

OpenStudy (anonymous):

you want \[y'\] and you know \[x'=26\] and the hint is that \[y^2=6^2+x^2+2\times 6\times x\cos(145)\]

OpenStudy (anonymous):

damn i meant \[y^2=6^2+x^2-12x\cos(145)\] not plus

OpenStudy (anonymous):

take the derivative, get \[2yy'=2xx'-12\cos(145)x'\] then plug in the numbers to get y'

OpenStudy (anonymous):

put \[x'=26\] \[x=3\times 26=78\] you still need y which you get from \[y^2=6^2+78^2-12\times 78(145)\]

OpenStudy (lgg23):

and \[y^{2} = 6^{2}+78^{2}\times78\cos (145)\] but thanks this really helped.